OpenStudy (anonymous):

If a resistance of R ohms is connected across a battery of E volts with internal resistance (the resistance within the battery) r ohms, then the power in watts dissipated in the external resistor is given by P = (E^2R)/((R+r)^2. If E and r are fixed, find R so that the power dissipated externally is a maximum.

OpenStudy (anonymous):

@Zarkon @mathstudent55 @Hero @iambatman

OpenStudy (anonymous):

It is an optimization problem for calculus...

OpenStudy (anonymous):

Have you studied implicit differentiation? This looks like it would be generally unpleasant to explicitly solve for R, then to differentiate. I would clear the fraction, implicitly differentiate the function, then set dR equal to zero. This should give you a relatively simple function of E and r.

OpenStudy (radar):

Or you could do a Google search of "Maximum Power Transfer Theorem"

OpenStudy (kropot72):

The power in the external resistance R is given by: \[\large P=E^{2} \times \frac{R}{(R+r)^{2}}\] \[\large \frac{dP}{dR}=E^{2} \times \frac{(R+r)^{2}-R.2(R+r)}{(R+r)^{4}}\] For a maximum or a minimum value of P, dP/dR = 0

OpenStudy (kropot72):

This happens when the numerator of the right hand side is zero, which occurs when: \[\large (R+r)^{2}=2R(R+r)\] This is satisfied by R = r. In order to demonstrate that this actually represents a maximum value for P, you must differentiate again and note the sign of d^2P/dR^2. Are you able to do this final step?

OpenStudy (anonymous):

I misread the question. kropot covered it well, though I thought we were supposed to guide to an answer, not solve the problem.

OpenStudy (kropot72):

@ZippySlug Thank you. Please note that I have left it for @ComputerNerd to finish by differentiating again to demonstrate that the tentative solution represents a maximum value for P.

OpenStudy (radar):

An excellent response kropot72, I thought a diagram and another idea may be helpful. Expanding dividing numerator and denominator by RL gettig: It is clear that The load resistor RL would equal the Internal resistance Rint for maximum transfer of power. Note that the efficiency is only 50% as you are only getting 1/2 the power generated into the load. If you are looking for maximum efficiency, matching RL to Ri isn't the way to go.